OpenStudy (anonymous):

Solve: (-2/(x+4))-(x+2)/(x-1)= (-3x)/(x^2+3x-4)

OpenStudy (anonymous):

Click "show steps" when you get there

OpenStudy (anonymous):

I must admit...wolfram alpha's steps are not the greatest

OpenStudy (anonymous):

once you get to the quadratic x^2 + 5x + 6 = 0, you can just write: (x+2)(x+3) = 0 x = -2, -3

OpenStudy (anonymous):

wow I feel so stupid, thank you so much :)